# Can Machines Resonate with Humans? Evaluating the Emotional and Empathic Comprehension of LMs

### v1

@@ 5 Collective Human Opinions
To analyze the subjectivity in empathic similarity labeling, we first collected eight annotations for each English story pairs, and then we investigated the impact of the language/culture on the empathy labeling by collecting and annotating Urdu story pairs.
@@ 5.1 English Story Pair Annotation
  • Continuous score on the summary: the annotators assign similarity scores from event, emotion, moral, and empathy as continuous value ranging from 1 to 4, based on the story summary.
@@
  • Discrete class-label on full story: considering that continuous scores provide larger labeling space, which may exacerbate subjectivity and inner-annotator disagreement, the annotators are asked to rate event and emotion similarity using class labels: very similar (V), moderately similar (M), and not similar (N). Here, we did not rate moral and empathy given they are too abstract concept to perceive and rate.
@@
Annotator Background
We had eight annotators, aged between 20 and 35 years, from diverse cultural and ethnic backgrounds, including two native Chinese speakers and six native Urdu speakers, with a balanced gender distribution. All annotators are proficient in English, some majored in psychology with bachelor’s degree and some are PhD and postdoc students in NLP. Training was performed before the formal annotations, which instructed the annotators how to rate the similarity score and highlighted the differences between event, emotion, moral, and empathy (see the annotation guidelines in Figure 5).
Results and Analysis
@@
This indicates that individuals having similar background and experiences tend to reach agreement in interpretations of the stories, particular for subtle and abstract concepts. Moreover, varying the agreement scores across different groups suggests that individual background and culture significantly influence how stories are perceived and annotated. This subjectivity poses challenges for models to learn and to replicate human judgements, especially for moral and empathy. The agreement of collective human opinions in similarity score annotations also presents an upper for a model.
@@
Given that none of the annotators is an English native speaker or has grown in a Western culture, bias might be introduced in labeling English stories pairs. We thus collected Urdu stories and asked Urdu native speakers to annotate them. This aims to eliminate potential biases, and most importantly to investigate the impact of languages and culture on the empathy similarity labeling.
@@ 5.2 Urdu Dataset Construction
Urdu is widely spoken in many South Asian countries, including Pakistan, India, Bangladesh, and Nepal. Roman Urdu/Hindi, which uses the English alphabet to write Urdu or Hindi, is commonly used in these regions for communication, discussions, and sharing feelings on media platforms. Consequently, we collected stories written in Roman Urdu to capture the authentic expression and nuances of this widely used form of communication.
@@
Synthetic Story Pairs Generation
@@
Human Annotations
Four Urdu native speakers were trained to annotate similarity scores for event, emotion and empathy, ranging from 1 to 4 with increments of 0.5; we excluded moral due to high ambiguity. The IAA between the four annotators is shown in Table 7. The last column overall represents the agreement between the averaged empathy scores of the four raters with the overall similarity score s provided by GPT-4o during story generation. Interestingly, empathy achieved the highest IAA, followed by emotion and then event, which differs with the findings in English data: event>emotion>empathy. This confirms highly subjective nature of the task. The correlations between the averaged empathy score and GPT-4o similarity are closer to those between annotators, suggesting the human-like judgement of GPT-4o.
LLM Estimations
@@ 6 Conclusion and Future Work
In future work, we will explore task reformulations to reduce the variability and subjectivity, and we will try more robust approaches to effectively model empathy in narratives.
@@
Appendix A Related Work
Emotion and Empathy in LMs
Many prior work focus on tasks of recognizing (mixed) emotions, scoring intensity, inferring and explaining person’s emotional reactions (Liu et al., 2024b; Chen et al., 2024). Sharma et al. (2020) collect EPITOME dataset to facilitate understanding empathy levels in online mental health conversations. Based on this dataset, Sharma et al. (2022) explore how human-AI collaboration could enhance empathic communication in text-based mental health support. They find that peer supporters can effectively use AI-generated feedback to improve their responses to support seekers. Sharma et al. (2021) further automate this process by leveraging a reinforcement learning agent, which makes sentence-level edits on supporters’ posts to increase empathy while maintaining conversation quality.
Recent study explored in general domains illustrates that current LLMs are far better than human in generation but fall short of understanding (West et al., 2023). However, for empathy-demanding scenarios, understanding and resonating with support-seekers are even more important than outputting opinions Buecker et al. (2021). In this work, we aim to examine language models capability in understanding and identifying nuanced difference of emotion and empathy, based on the task of empathic similarity (Shen et al., 2023). It examines models understanding of human narratives in both emotion and empathy by assessing the similarity between two stories from event, emotion, moral and empathy. However, predictions show low correction with human-annotated labels by both fine-tuning LMs (e.g., SBERT and BART) on empathic similarity dataset and prompting GPT-3.5-turbo. We speculate that the sentence representations may capture adequate semantic context, but lack emotional and empathic signals. In this work, we aim to improve the estimation accuracy by forcing models to discern and learn the underlying reasons for similarities between pairs of stories, via (1) enhancing LM-based sentence embeddings, and (2) empowering LLM reasoning capability.
Sentence Embedding Enhancement
Many previous methods improve semantic embeddings by adjusting loss functions in training. SimCSE leverages contrastive loss Gao et al. (2021), and ESimCSE applies momentum contrast strategy to increase the number of negative pairs involved in the loss calculation, showing notable improvements across multiple semantic textual similarity benchmarks (Wu et al., 2021). Hubert et al. (2024) improved the knowledge graph embeddings by integrating semantic awareness into traditional loss functions like pairwise hinge loss, binary cross-entropy loss, and logistic loss. Huang et al. (2024) use a multi-task hybrid loss, incorporating both skeletal and semantic embedding into loss functions for micro-gesture classification. Therefore, we investigate a variety of loss functions such as ContrastiveLoss, CosineSimilarityLoss, CoSENTLoss and AnglELoss, based on LMs including MPNet Song et al. (2020), RoBERTa (base, large) Liu et al. (2019), DeBERTa (small, large) He et al. (2020) to improve sentence embedding in representing empathic features.
LLM Reasoning and Fine-tuning
Naive LLM prompts may fail to fully leverage LLMs’ reasoning capabilities, leading to poor accuracy on EmpathicStories dataset Shen et al. (2023). Recent advancements in zero- and few-shot prompting techniques may boost performance such as Chain-of-Thought prompting (CoT) Wei et al. (2022), Least-to-Most prompting Zhou et al. (2022), and search-based approaches like Tree-of-Thought (ToT) Yao et al. (2024). We prompt LLMs by CoT and also fine-tune LLMs with reasoning trajectories, encouraging LLMs to think over before making the final decision.
Appendix B Error Analysis
@@
Appendix C Experimental Results
C.1 Discriminative Models Results
@@
C.2 Generative LLMs Results
@@
Table 12: Results of Generative LLMs under zero-shot configuration using the full and summary version story. Classification gold labels are binned by $\text{score}>2.5$.
Figure 4: Confusion matrix of the first digit prediction.
Appendix D Annotation Guidelines
@@
Appendix E Urdu Story Pair Generation
@@
Appendix F English Story Pair Example
This figure shows a sample pair of stories along with their summarized versions and scores in four aspects: empathy, event, emotion, and moral.

### v2

@@ 6 Collective Human Opinions
To analyze the subjectivity in empathic similarity labeling, we first collected eight annotations for each pair of English stories. We then explored the impact of language and culture on empathy labeling by collecting and annotating Urdu story pairs.
@@ 6.1 English Story Pair Annotation
  • Continuous score on the summary: The annotators assigned similarity scores for event, emotion, moral, and empathy as values ranging from 1 to 4, based on the story summaries.
@@
  • Discrete class-label on full story: Considering that continuous scores provide larger labeling space, which may exacerbate subjectivity and inner-annotator disagreement, the annotators are asked to rate event and emotion similarity using class labels: very similar (V), moderately similar (M), and not similar (N). Here, we did not rate moral and empathy given that the concepts were too abstract to perceive and rate.
@@
Annotator Background:
We had eight annotators, aged between 20 and 35 years, from diverse cultural and ethnic backgrounds, including two native Chinese speakers and six native Urdu speakers, with a balanced gender distribution. All annotators are proficient in English, some majored in psychology with bachelor’s degree and some are PhD and postdoc students in NLP. Training was performed before the formal annotations, which instructed the annotators how to rate the similarity score and highlighted the differences between event, emotion, moral, and empathy (guidelines in Figure 6).
Results and Analysis:
@@
This indicates that individuals who have closer interpersonal relationships with each other reach better agreement in interpretations of the stories, particular for subtle and abstract concepts. Moreover, varying the agreement scores across different groups suggests that individual background and culture significantly influence how stories are perceived and annotated. This subjectivity poses challenges for models to learn and to replicate human judgements, especially for moral and empathy. The agreement of collective human opinions in similarity score annotations also presents an upper bound for a model.
@@ 6.2 Urdu Dataset Construction
Given that none of the annotators are native English speakers nor have grown up in a Western culture, bias might be introduced in labeling English stories pairs. We thus collected Urdu stories and asked Urdu native speakers to annotate them. This aims to eliminate potential biases, and most importantly to investigate the impact of languages and culture on the empathy similarity labeling. Urdu is widely spoken in many South Asian countries, including Pakistan, India, Bangladesh, and Nepal. Roman Urdu/Hindi, which uses the English alphabet to write Urdu or Hindi, is commonly used in these regions for communication, discussions, and sharing feelings on media platforms. Consequently, we collected stories written in Roman Urdu to capture the authentic expression and nuances of this widely used form of communication.
@@
Synthetic Story Pairs Generation:
@@
Human Annotations:
Four Urdu native speakers were trained to annotate similarity scores for event, emotion and empathy, ranging from 1 to 4 with increments of 0.5; we excluded moral due to high ambiguity. The IAA between the four annotators is shown in Table 7. The last column overall represents the agreement between the averaged empathy scores of the four raters with the overall similarity score s provided by GPT-4o during story generation. Interestingly, empathy achieved the highest IAA, followed by emotion and then event, which differs with the findings in English data: event>emotion>empathy.
Agreement scores can be significantly influenced by shared interpersonal traits among annotators. Cultural factors also play a role, with closer interpersonal relationships leading to higher agreement scores. As demonstrated in the results, when annotators are more closely related, resonate with each other and have the same perception about the world, their agreement score is higher which may enhance the LMs capacity to learn cultural specific empathy. This suggests that, enhancing the machines’ understanding of empathy, stories must be collected and annotated within targeted countries or regions where there is a deep understanding of local demographic norms. The correlations between the averaged empathy score and GPT-4o similarity are closer to those between annotators, suggesting the human-like judgement of GPT-4o.
LLM Estimations:
@@ 7 Conclusion and Future Work
We find that many factors impact the inner annotator agreement in empathic similarity labeling, including the native language of annotators, the similarity of their background, experience, the closeness between them, and training process. Considering these factors that affect the correlations between annotators in a unified framework would be an interesting direction to explore in future work, especially the studies across multiple languages. Furthermore, we will explore task reformulations to reduce the variability and subjectivity, and we will try more robust approaches to effectively model empathy in narratives.
@@
Appendix A Error Analysis
Figure 2 shows the correlation between the four similarity scores. Table 9 exhibits the results authors mentioned in Shen et al. (2023) and our reproduction results on EmpathicStories. Figure 3 show the prediction count with SBERT and BART on test and development set.
@@ Appendix A Error Analysis
Appendix B Experiments
B.1 Gold-label Guided Explanation Generation
We used Llama3-70B-instruct with the prompts detailed in Figure 4 to analyze the story pairs across various dimensions, generating analysis that was subsequently used as reasoning content for supervised fine-tuning (SFT).
Figure 4: Gold-label Guided Explanation Generation Prompt using Llama3-70B-instruct.
B.2 Discriminative Models Results
@@
B.3 Generative LLMs Results
Zero-shot Generative LLMs.
Table 3 demonstrates the results of zero-shot generative LLMs including Llama-3-8B, Llama-3-70B and GPT-4o on test set, Table 12 provides the additional results over development set.
@@ B.3 Generative LLMs Results
Table 12: Zero-shot Generative LLMs results using the full story vs. story summary, over four types of gold similarity scores: empathy, event, emotion and moral. Classification gold labels are binned by score > 2.5.
Understanding Bottleneck of Fine-tuned LLMs.
Figure 5 shows the confusion matrix predicted by fine-tuned Llama-3-8B on training set. The model learned nothing but statistical $P(Y)$ of training set, leading to random predictions conditioned on input story pairs. See more in Section 5.3.
Figure 5: Confusion matrix of fine-tuned Llama-3-8B on training set. The model could not estimate the corresponding score conditioned on the input story pair, but sampled a similarity class based on the gold class distribution of training data $P(Y)$ = (0.140, 0.399, 0.404, 0.057) whatever the input pair was, leading to randomness on seen and unseen cases prediction. It learned nothing but statistical $P(Y)$ of training set. See more in Section 5.3.
Appendix C Annotation Guidelines
@@
Appendix D Urdu Story Pair Generation
@@
Appendix E English Story Pair Example
The illustrative example form the EmpathicStories dataset Shen et al. (2023) shows a sample pair of stories along with their summarized versions and scores in four aspects: empathy, event, emotion, and moral.